The cross product of two unit vector gives rise to -------?


1
Expert's answer
2021-03-21T11:36:47-0400

"\vec i \times\vec j=\vec k"


"\vec j \times\vec i=-\vec k"


"\vec j \times\vec k=\vec i"


"\vec k \times\vec j=-\vec i"


"\vec k \times\vec i=\vec j"


"\vec i \times\vec k=-\vec j"


"\vec i \times\vec i=\vec j \times\vec j=\vec k \times\vec k=0"